OpenStudy (anonymous):

seth can read 6 3/4 books in 2 1/4 weeks how many books can seth read in one week. A. 3 2/3 b. 1/3 c. 15 8/9 d. 3

OpenStudy (anonymous):

(6 3/4)/(2 1/4) = (27/4)/(9/4) = 3 books/week

OpenStudy (anonymous):

Change the mixed numbers to improper fractions: 6 3/4 = ((4times6)+3)/4 = 27/4 2 1/4 = ((4times2)+1)/4 = 9/4 Set up a proportion: 27/4 books / 9/4 weeks = x books / 1 week (27/4) / (9/4) = x / 1 (27/4) / (9/4) = x (27/4) times (4/9) = x (27times4)/(4times9) = x 3 = x Answer: 3 books *Of course, you could have also divided 27/4 by 9/4 from the very beginning. However, it's good to know how to solve using proportions for those times like if this problem had been anything other than 1 week.